OpenStudy (anonymous):

Develop a formula Cosine (4 theta) as a fourth-degree polynomial in the variable equals what cosine (theta)?

OpenStudy (asnaseer):

have you learnt about \(e^i\theta=\cos(\theta)+i\sin(\theta)\)?

OpenStudy (asnaseer):

sorry that should be \(e^{i\theta}\)

OpenStudy (anonymous):

I think he/she want taylor approximation

OpenStudy (asnaseer):

why do you say that?

OpenStudy (anonymous):

something about fourth degree polynomial

OpenStudy (asnaseer):

well it will involve terms like \(\cos^4(\theta)\)

OpenStudy (asnaseer):

maybe @coolcat93 can confirm for us...